# Iker García-Ferrero, Begoña Altuna HiTZ Basque Center for Language Technology - Ixa

### v1

@@ 5.1 Zero-Shot summarization
The results indicate that GPT-4 generates the highest quality summaries, whereas Nous-Hermes-2-Yi-34B produces the shortest summaries without compromising quality. Both, Nous-Hermes-2-SOLAR-10.7B and OpenChat-3.5-0106, with 10.7 billion and 7 billion parameters respectively, deliver remarkably good performance despite being significantly smaller than the 34 billion parameter model, Nous-Hermes-2-Yi-34B. This makes these two models the preferred options when considering computational resource constraints.
The zero-shot results indicate that for this task, the number of model parameters holds little significance compared to the pre-training data. For instance, Llama-2-70b-chat-hf, a 70 billion parameter model, underperforms compared to the OpenChat-3.5-0106 model, which has only 7 billion parameters. The models that consistently achieve better performance share a common feature: they are trained on very large high-quality synthetic datasets. For example, the Nous-Hermes series of models have been trained using the OpenHermes dataset [Teknium,2023], which comprises 1 million instructions generated mainly with GPT-4. Similarly, OpenChat-3.5-0106 was also trained using data generated by GPT-4 and GPT-3.5 [Wang et al.,2023]. Deepseek [Bi et al.,2024] was trained with 1.5 million instructions. Further analysis of the publicly available OpenHermes 2.5 dataset through the Lilac platform  [^9] https://lilacai-lilac.hf.space/datasets#lilac/OpenHermes-2.5 reveals that this dataset contains at least 25,749 examples of summarization, a significant portion of which are article summarizations. Therefore, this suggests that the best-performing models in our dataset are those that have been exposed to a large amount of summarization examples during fine-tuning. It’s important to note that the OpenHermes dataset predominantly contains instructions in English, with only 62 of the 1 million instructions being in Spanish, and these are specific examples of translations. This highlights the models’ remarkable proficiency in transfer learning, demonstrating their ability to effectively apply knowledge learned from English data to tasks in Spanish.
5.2 Error analysis
To better understand the models’ performance on our dataset, we conducted an error analysis to identify the circumstances and reasons why the models failed. We identified three different error sources:
@@ 5.2.1 Evaluation Consistency with Different Human Annotations
Figure 5: ROUGE scores of the models when evaluated against the gold summaries and the validation summaries produced by a second set of human annotators.
@@
As an example, in Table 6, we analyze the performance of Nous-Hermes-2-SOLAR-10.7B when trained with varying amounts of training samples. The results indicate that when trained with a small number of samples, the model can learn to produce concise summaries, reducing the average number of words in the output summary by 2.5. On the other hand, the ROUGE score continues to increase as more training samples are added. The model’s performance does not show any signs of saturation. This suggests that extending the dataset beyond 750 training samples could further enhance model performance.
@@ Acknowledgements
A demo of the ClickbaitFighter-7B model can be found in the SomosNLP Hugging Face organization. This demo was developed during the SomosNLP 2024 Hackathon: https://hf.co/spaces/somosnlp/NoticIA-demo
@@ Appendix A Fine tuning details
In this section, we detail our fine-tuning process. We fine-tune all model parameters using bfloat16 precision. Utilizing DeepSpeed Zero 3 [Ren et al.,2021], we distribute the model, gradients, and optimizer across four A100 80GB GPUs. We employ the standard Next Token Prediction (NTP) loss for training our models. To prevent the loss associated with the article body tokens from overshadowing the loss of the summary output tokens, we compute the loss exclusively over the summary tokens. The hyperparameter settings are outlined in Table 6. For all the experiments, we use greedy search to generate text.
ClickbaitFighter-2B ClickbaitFighter-10B Batch Size 64 64 Optimizer AdamW AdamW Scheduler Cosine Cosine Learning Rate 0.000005 0.000005 Weight Decay 0.0 0.0 Warmup Ratio 0.1 0.1 adam_beta1 0.9 0.9 adam_beta2 0.95 0.95 adam_epsilon 1e-12 1e-12 Epochs 5 3 Sequence Length 8192 8192 Compute Type Bfloat16 Bfloat16
@@ Appendix D Model details
Model Family Model name Parameter Count Citation URL Deepseek deepseek-llm-67b-chat 67B [Bi et al.,2024] https://hf.co/deepseek-ai/deepseek-llm-67b-chat Gemma gemma-2b-it 2.51B [Gemma-Team et al.,2024] https://hf.co/google/gemma-2b-it gemma-7b-it 8.54B [Gemma-Team et al.,2024] https://hf.co/google/gemma-7b-it Llama2 Llama-2-70b-chat-hf 70B [Touvron et al.,2023] https://hf.co/meta-llama/Llama-2-70b-chat-hf Llama-2-13b-chat-hf 13B [Touvron et al.,2023] https://hf.co/meta-llama/Llama-2-13b-chat-hf Llama-2-7b-chat-hf 7B [Touvron et al.,2023] https://hf.co/meta-llama/Llama-2-7b-chat-hf Nous-Hermes-2-Llama-2-70B.yaml 70B [Teknium,2023] https://hf.co/NousResearch/Nous-Hermes-2-Llama-2-70B TinyLlama-1.1B-Chat-v1.0 1.1B [Zhang et al.,2024] https://hf.co/TinyLlama/TinyLlama-1.1B-Chat-v1.0 tulu-2-dpo-70b 70B [Ivison et al.,2023] https://huggingface.co/allenai/tulu-2-dpo-70b Mistral Mistral-7B-Instruct-v0.2 7B [Jiang et al.,2023] https://hf.co/mistralai/Mistral-7B-Instruct-v0.2 Nous-Hermes-2-Mistral-7B-DPO 7B [Teknium,2023] https://hf.co/NousResearch/Nous-Hermes-2-Mistral-7B-DPO Nous-Hermes-2-Pro-Mistral-7B 7B [Teknium,2023] https://hf.co/NousResearch/Hermes-2-Pro-Mistral-7B openchat-3.5-0106 7B [Wang et al.,2023] https://hf.co/openchat/openchat-3.5-0106 OpenHermes-2.5-Mistral-7B 7B [Teknium,2023] https://hf.co/teknium/OpenHermes-2.5-Mistral-7B zephyr-7b-beta 7B [Tunstall et al.,2023] https://hf.co/HuggingFaceH4/zephyr-7b-beta Solar SOLAR-10.7B-Instruct-v1.0 10.7B [Kim et al.,2023] https://hf.co/upstage/SOLAR-10.7B-Instruct-v1.0 Nous-Hermes-2-SOLAR-10.7B 10.7B [Teknium,2023] https://hf.co/NousResearch/Nous-Hermes-2-SOLAR-10.7B Mixtral Mixtral-8x7B-Instruct-v0.1 46.7B [Jiang et al.,2024] https://hf.co/mistralai/Mixtral-8x7B-Instruct-v0.1 Nous-Hermes-2-Mixtral-8x7B-DPO 46.7B [Teknium,2023] https://hf.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Nous-Hermes-2-Mixtral-8x7B-SFT 46.7B [Teknium,2023] https://hf.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-SFT StableLM Nous-Capybara-3B-V1.9 2.8B [Teknium,2023] https://hf.co/NousResearch/Nous-Capybara-3B-V1.9 rocket-3B 2.8B [pansophic,2023] https://hf.co/pansophic/rocket-3B stablelm-zephyr-3b 3B [stability.ai,2023] https://hf.co/stabilityai/stablelm-zephyr-3b Yi Nous-Hermes-2-Yi-34B 34B [Teknium,2023] https://hf.co/NousResearch/Nous-Hermes-2-Yi-34B Yi-6B-Chat 6B [01.AI et al.,2024] https://hf.co/01-ai/Yi-6B-Chat Yi-34B-Chat 34B [01.AI et al.,2024] https://hf.co/01-ai/Yi-34B-Chat QWEN Qwen1.5-7B-Chat 7B [Bai et al.,2023] https://hf.co/Qwen/Qwen1.5-7B-Chat Qwen1.5-14B-Chat 14B [Bai et al.,2023] https://hf.co/Qwen/Qwen1.5-14B-Chat Qwen1.5-72B-Chat 72B [Bai et al.,2023] https://hf.co/Qwen/Qwen1.5-72B-Chat Smaug-72B-v0.1 72B [Pal et al.,2024] https://hf.co/abacusai/Smaug-72B-v0.1 ChatGPT gpt-4-0125-preview unknown [OpenAI,2023] https://openai.com/ gpt-3.5-turbo-0125 unknown [OpenAI,2024] https://openai.com/
@@ Appendix E Extended Results
In this section, we present all the results in numerical format. The results can be found in Table 8. Additionally, we include models that, due to their significantly low performance, were not featured in the main results section.
@@ Appendix F NoticIA Dataset: Annotator Guidelines
Model Name ROUGE Score Average Summary Length Human 56.43 11.03 ClickbaitFighter-10B 52.01 7.66 ClickbaitFighter-7B 49.81 8.48 ClickbaitFighter-2B 37.58 9.23 gpt-4-0125-preview 32.00 23.82 Nous-Hermes-2-Yi-34B 29.99 18.98 deepseek-llm-67b-chat 29.09 25.15 gpt-3.5-turbo-0125 28.64 20.44 openchat-3.5-0106 28.42 23.61 Qwen1.5-72B-Chat 27.80 26.89 Nous-Hermes-2-SOLAR-10.7B 27.48 20.72 Nous-Hermes-2-Mixtral-8x7B-SFT 26.99 31.89 Nous-Hermes-2-Mistral-7B-DPO 26.33 25.22 OpenHermes-2.5-Mistral-7B 25.91 25.21 Mistral-7B-Instruct-v0.2 25.31 26.01 Mixtral-8x7B-Instruct-v0.1 25.24 32.39 Qwen1.5-14B-Chat 23.88 25.57 tulu-2-dpo-70b 23.27 38.56 gemma-2b-it 22.24 24.61 gemma-7b-it 22.04 28.19 SOLAR-10.7B-Instruct-v1.0 21.34 38.13 Qwen1.5-7B-Chat 20.66 26.26 Nous-Hermes-2-Mixtral-8x7B-DPO 19.74 58.68 Llama-2-70b-chat-hf 19.20 59.15 Llama-2-13b-chat-hf 18.37 54.46 zephyr-7b-beta 15.64 105.75 Llama-2-7b-chat-hf 14.65 73.50 stablelm-zephyr-3b 12.02 86.90 Nous-Hermes-2-Llama-2-70B 7.22 154.55 TinyLlama-1.1B-Chat-v1.0 3.84 191.15 Smaug-72B-v0.1 1.79 82.74 rocket-3B 0.92 228.43

### v2

@@ 5.1 Zero-Shot summarization
The results indicate that Llama-3-70B-Instruct generates the highest quality summaries, with GPT-4o and GPT-4 following closely. Both Nous-Hermes-2-SOLAR-10.7B and OpenChat-3.5-0106, with 10.7 billion and 7 billion parameters respectively, deliver remarkably good performance despite being significantly smaller than the 70 billion parameter Llama-3 model. This makes these two models preferred the options when considering computational resource constraints.
The zero-shot results indicate that for this task, the number of model parameters holds little significance compared to the pre-training data. For instance, Llama-2-70b-chat-hf, a 70 billion parameter model, underperforms compared to the OpenChat-3.5-0106 model, which has only 7 billion parameters. The models that consistently achieve better performance share a common feature: they are trained on very large high-quality datasets. For example, the Nous-Hermes series of models have been trained using the OpenHermes dataset ([Teknium, 2023], which comprises 1 million instructions generated mainly with GPT-4. Similarly, OpenChat-3.5-0106 was also trained using data generated by GPT-4 and GPT-3.5 ([Wang et al., 2023]. Deepseek ([Bi et al., 2024] was trained with 1.5 million instructions. Further analysis of the publicly available OpenHermes 2.5 dataset through the Lilac platform [^10] https://lilacai-lilac.hf.space/datasets#lilac/OpenHermes-2.5 reveals that this dataset contains at least 25,749 examples of summarization, a significant portion of which are article summaries. Therefore, this suggests that the best-performing models in our dataset are those that have been exposed to a large amount of summary examples during fine-tuning. It’s important to note that the OpenHermes dataset predominantly contains instructions in English, with only 62 of the 1 million instructions being in Spanish, and these are specific examples of translations. This highlights the models’ remarkable proficiency in transfer learning, demonstrating their ability to effectively apply knowledge learned from English data to tasks in Spanish.
5.2 Error analysis
To better understand the models’ performance on our dataset, we conducted an error analysis to identify the circumstances and reasons why the models failed. We identified three different error sources for which we provide examples in Table 5:
@@ 5.2.1 Evaluation Consistency with Different Human Annotations
Figure 5: ROUGE scores of the models when evaluated against the gold summaries and the validation summaries produced by the second annotator.
@@
In Figure 6, we analyze the performance of Nous-Hermes-2-SOLAR-10.7B when trained with varying amounts of training samples. The results indicate that when trained with a small number of samples, the model can learn to produce concise summaries, reducing the average number of words in the output summary by 2.5. On the other hand, the ROUGE score continues to increase as more training samples are added and the model’s performance does not show any signs of saturation. This suggests that extending the dataset beyond 750 training samples could further enhance model performance.
@@ Acknowledgements
A demo of the ClickbaitFighter-7B model can be found in the SomosNLP Hugging Face organization. This demo was developed during the SomosNLP 2024 Hackathon [^11] https://somosnlp.org/ #Somos600M: https://hf.co/spaces/somosnlp/NoticIA-demo
@@ Appendix A Fine tuning details
We fine-tune all model parameters using bfloat16 precision. Utilizing DeepSpeed Zero 3 ([Ren et al., 2021], we distribute the model, gradients, and optimizer across four A100 80GB GPUs. We employ the standard Next Token Prediction (NTP) loss for training our models. To prevent the loss associated with the article body tokens from overshadowing the loss of the summary output tokens, we compute the loss exclusively over the summary tokens. The hyperparameter settings are outlined in Table 6. For all the experiments, we use greedy search to generate text.
ClickbaitFighter-2B ClickbaitFighter-7B/10B Batch Size 64 64 Optimizer AdamW AdamW Scheduler Cosine Cosine Learning Rate 0.000005 0.000005 Weight Decay 0.0 0.0 Warmup Ratio 0.1 0.1 adam_beta1 0.9 0.9 adam_beta2 0.95 0.95 adam_epsilon 1e-12 1e-12 Epochs 5 3 Sequence Length 8192 8192 Compute Type Bfloat16 Bfloat16
@@ Appendix D Model details
Model Family Model name Parameter Count Citation URL Deepseek deepseek-llm-67b-chat 67B ([Bi et al., 2024] https://hf.co/deepseek-ai/deepseek-llm-67b-chat Gemma gemma-2b-it 2.51B ([Gemma-Team et al., 2024] https://hf.co/google/gemma-2b-it gemma-7b-it 8.54B ([Gemma-Team et al., 2024] https://hf.co/google/gemma-7b-it Llama2 Llama-2-70b-chat-hf 70B ([Touvron et al., 2023] https://hf.co/meta-llama/Llama-2-70b-chat-hf Llama-2-13b-chat-hf 13B ([Touvron et al., 2023] https://hf.co/meta-llama/Llama-2-13b-chat-hf Llama-2-7b-chat-hf 7B ([Touvron et al., 2023] https://hf.co/meta-llama/Llama-2-7b-chat-hf Nous-Hermes-2-Llama-2-70B.yaml 70B ([Teknium, 2023] https://hf.co/NousResearch/Nous-Hermes-2-Llama-2-70B TinyLlama-1.1B-Chat-v1.0 1.1B ([Zhang et al., 2024] https://hf.co/TinyLlama/TinyLlama-1.1B-Chat-v1.0 tulu-2-dpo-70b 70B ([Ivison et al., 2023] https://hf.co/allenai/tulu-2-dpo-70b Llama3 Meta-Llama-3-8B-Instruct 70B ([AI@Meta, 2024] https://hf.co/meta-llama/Meta-Llama-3-8B-Instruct Meta-Llama-3-8B-Instruct 8B ([AI@Meta, 2024] https://hf.co/meta-llama/Meta-Llama-3-70B-Instruct Mistral Mistral-7B-Instruct-v0.2 7B ([Jiang et al., 2023] https://hf.co/mistralai/Mistral-7B-Instruct-v0.2 Nous-Hermes-2-Mistral-7B-DPO 7B ([Teknium, 2023] https://hf.co/NousResearch/Nous-Hermes-2-Mistral-7B-DPO Nous-Hermes-2-Pro-Mistral-7B 7B ([Teknium, 2023] https://hf.co/NousResearch/Hermes-2-Pro-Mistral-7B openchat-3.5-0106 7B ([Wang et al., 2023] https://hf.co/openchat/openchat-3.5-0106 OpenHermes-2.5-Mistral-7B 7B ([Teknium, 2023] https://hf.co/teknium/OpenHermes-2.5-Mistral-7B zephyr-7b-beta 7B ([Tunstall et al., 2023] https://hf.co/HuggingFaceH4/zephyr-7b-beta Solar SOLAR-10.7B-Instruct-v1.0 10.7B ([Kim et al., 2023] https://hf.co/upstage/SOLAR-10.7B-Instruct-v1.0 Nous-Hermes-2-SOLAR-10.7B 10.7B ([Teknium, 2023] https://hf.co/NousResearch/Nous-Hermes-2-SOLAR-10.7B Mixtral Mixtral-8x7B-Instruct-v0.1 46.7B ([Jiang et al., 2024] https://hf.co/mistralai/Mixtral-8x7B-Instruct-v0.1 Nous-Hermes-2-Mixtral-8x7B-DPO 46.7B ([Teknium, 2023] https://hf.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-DPO Nous-Hermes-2-Mixtral-8x7B-SFT 46.7B ([Teknium, 2023] https://hf.co/NousResearch/Nous-Hermes-2-Mixtral-8x7B-SFT StableLM Nous-Capybara-3B-V1.9 2.8B ([Teknium, 2023] https://hf.co/NousResearch/Nous-Capybara-3B-V1.9 rocket-3B 2.8B ([pansophic, 2023] https://hf.co/pansophic/rocket-3B stablelm-zephyr-3b 3B ([stability.ai, 2023] https://hf.co/stabilityai/stablelm-zephyr-3b Yi Nous-Hermes-2-Yi-34B 34B ([Teknium, 2023] https://hf.co/NousResearch/Nous-Hermes-2-Yi-34B Yi-6B-Chat 6B ([01.AI et al., 2024] https://hf.co/01-ai/Yi-6B-Chat Yi-34B-Chat 34B ([01.AI et al., 2024] https://hf.co/01-ai/Yi-34B-Chat QWEN Qwen1.5-7B-Chat 7B ([Bai et al., 2023] https://hf.co/Qwen/Qwen1.5-7B-Chat Qwen1.5-14B-Chat 14B ([Bai et al., 2023] https://hf.co/Qwen/Qwen1.5-14B-Chat Qwen1.5-72B-Chat 72B ([Bai et al., 2023] https://hf.co/Qwen/Qwen1.5-72B-Chat Smaug-72B-v0.1 72B ([Pal et al., 2024] https://hf.co/abacusai/Smaug-72B-v0.1 ChatGPT gpt-4o-2024-05-13 unknown ([OpenAI, 2024b] https://openai.com/ gpt-4-turbo-2024-04-09 unknown ([OpenAI, 2023] https://openai.com/ gpt-3.5-turbo-0125 unknown ([OpenAI, 2024a] https://openai.com/
@@ Appendix E Extended Results
Table 8 displays the results for the assessed models in numerical format. Additionally, we include models that, due to their significantly low performance, were not featured in the main results section.
@@ Appendix F NoticIA Dataset: Annotator Guidelines
Model Name ROUGE Score Average Summary Length Human 55.98 11.03 ClickbaitFighter-10B 52.01 7.66 ClickbaitFighter-7B 49.81 8.48 Meta-Llama-3-70B-Instruct 36.30 20.72 ClickbaitFighter-2B 36.26 8.04 gpt-4o 35.45 21.81 gpt-4-turbo 34.03 23.86 gpt-4-0125-preview 32.00 23.82 Nous-Hermes-2-Yi-34B 29.99 18.98 deepseek-llm-67b-chat 29.09 25.15 openchat-3.5-0106 28.42 23.61 Qwen1.5-72B-Chat 27.80 26.89 Nous-Hermes-2-SOLAR-10.7B 27.48 20.72 Meta-Llama-3-8B-Instruct 27.21 30.72 gpt-3.5-turbo-0125 27.08 20.72 Nous-Hermes-2-Mixtral-8x7B-SFT 26.99 31.89 Nous-Hermes-2-Mistral-7B-DPO 26.33 25.22 OpenHermes-2.5-Mistral-7B 25.91 25.21 Mistral-7B-Instruct-v0.2 25.31 26.01 Mixtral-8x7B-Instruct-v0.1 25.24 32.39 Qwen1.5-14B-Chat 23.88 25.57 Hermes-2-Pro-Mistral-7B 23.50 47.77 tulu-2-dpo-70b 23.27 38.56 gemma-2b-it 22.24 24.61 gemma-7b-it 22.04 28.19 SOLAR-10.7B-Instruct-v1.0 21.34 38.13 Qwen1.5-7B-Chat 20.66 26.26 Nous-Hermes-2-Mixtral-8x7B-DPO 19.74 58.68 Llama-2-70b-chat-hf 19.20 59.15 Llama-2-13b-chat-hf 18.37 54.46 zephyr-7b-beta 15.64 105.75 Llama-2-7b-chat-hf 14.65 73.50 stablelm-zephyr-3b 12.02 86.90 Nous-Hermes-2-Llama-2-70B 7.22 154.55 TinyLlama-1.1B-Chat-v1.0 3.84 191.15 Smaug-72B-v0.1 1.79 82.74 rocket-3B 0.92 228.43